OpenStudy (anonymous):

Find the number of solutions of each equation using the discriminant. x^2+3x-4=0

OpenStudy (amistre64):

b^2 -4ac is the discriminate

OpenStudy (anonymous):

how did you get the answer

OpenStudy (amistre64):

3^2 -4(1)(-4) = 9+16 = graeter than 0; 2 real solutions

OpenStudy (amistre64):

there are 2 parts to the quadratic formula; the determinate and the discriminate; the determinate determines the place you stand in the middle..... the discriminate tells you where to look for solutions from that point :)

OpenStudy (amistre64):

if the discriminate is zero; the only solution is where you are at. 1 solution if the discriminate is positive; then there are 2 solutions at equal distances to the left of you, and one to the right of you

OpenStudy (amistre64):

if the discriminate is a (-) number; that means the solutions are above you and are not part of where you are standing at all.... no solutions where you are :)

OpenStudy (amistre64):